what is the magnitude
U: magnitude 23, bearing 215 degrees
V: magnitude 14.5, bearing 105 degrees
W: U+V

Answer by venugopalramana(3286)

U: magnitude 23, bearing 215 degrees=23(COS(215)+iSIN(215))=-18.866-13.156i
V: magnitude 14.5, bearing 105 degrees=14.5(COS(105)+iSIN(105)=-3.74+14.01i
W: U+V=-22.606+0.854i
|W|=SQRT(22.606^2+0.854^2)=22.622